NAOJ GW Elog Logbook 3.2

From 23-12-26
Yuhang, Michael
Apparently there was some issue with the connection of rampeauto to main laser (filter cavity servo). This was turned off and CC PLL improved, but the issue is still delayed into the future.
The current situation with the CC PLL is that the cable from the slow loop correction of the servo to the CC laser wasn't working, so we improvised and it's dangling over the optical table. This will need to be fixed at some point. The issue with the 2V required/200 mV provided correction signal was due to the sideband being on the wrong side of resonance. We should be such that decreaing the temperature of the laser decreases the frequency of the peak on the spectrum analyzer, and then position the peak about 20-30 MHz away from resonance before turning on the loop. This is a strange issue of the CC PLL system where somehow the correction signal has some offset and only properly corrects for sidebands initially far enough away from resonance.
We wanted to test taking the data for the Taiwan machine learning measurement. We injected to CC2 Ramp IN but it showed little if no effect on homodyne. In the electrical diagram EPS1 and Perturb IN add to give EPS2 so we just decided to add it at Perturb IN. Then we can see what looks like squeezing and antisqueezing on the homodyne. However, the maximum voltage sent to Perturb IN while CC2 is locked is quite low, only up to about 10 mV. Still, it seems to work, so we should verify squeezing and antisqueezing level over the scan range using the DGS system.

From 23-12-26
Yuhang, Hsien-Yi, Michael
We discussed taking data for the Taiwan group machine learning project. We would like to take a long scan of homodyne subtraction DC port while continuously scanning the homodyne angle, in order to have an appropriate time series for FIS training data. The proposition is to scan the phase of the CC2 servo, which controls the phase of the local oscillator via the infrared phase shifter, while CC2 is locked, and then leave it for some time with several periods of squeezing/antisqueezing traversal. Typically we set homodyne angle in DDS through DDS3 DAC3 CC2 Demodulation (though it's actually Channel 1 in the software). So we figured to inject a ramp signal to CC2 Perturb IN, which in principle should allow us to scan the homodyne angle. Unfortunately homodyne is demodulated using a single chanel mixer. We should buy an IQ mixer to allow us to tell if it is amplification or deamplification.

Yuhang, Michael
We investigated the cause of the CC PLL glitches and saw that it was not so stable as it seemed on Friday. I had to leave earlier but Yuhang continued to investigate afterward (next entry).
On the spectrum analyzer we normally look at the DC peak and visualize the CC PLL sidebands at -7 and 7 MHz. Whenever the system unlocks it seems to be precluded by small bumps about 1 MHz away from the sidebands.
Another thing to check was the magnitude of the correction signal. The CC laser has a piezo tuning coefficient of ~ 1 V/MHz. If we move the sidebands about 2 MHz away from lock, we see that the fast loop monitor output only shows 200 mV, which isn't enough to properly correct, though it moves in the correct direction. From the specifications of the ADF4001 board the fast loop should be able to provide 10 V. It seems that there is definitely some problem with the electronics. We see that ppol PLL gives roughly the correct magnitude of correction signal (2 V for 3 MHz detuning from 160 MHz).
We once again thought it was good enough, and decided to move on to characterizing optical loss and phase noise by measuring the sqz/asqz ratio for different green pump powers, since the purpose of switching to the "one EOM" layout was to reduce squeezing phase noise.
However, we then ended up seeing some CC2 phase jitter despite the board saying it was locked. We checked CC1 and CC2 error signals and saw that the glitches were correlated in each of these channels. At this point we were not sure whether or not the glitches were coming from the lasers or the electronics.
We sent the PLL CC monitor channel to a mixer and demodulated at 7 MHz. It should be flat but instead we saw spikes. We tried also for ppol, reducing the modulation to 80 MHz (DDS has a bandwidth of 110 MHz), and then demodulated at 80 MHz. We once again saw glitches in the ppol. Looking at the time domain SHG error signal on lock shows none of these glitches though. If the problem is in the main laser, then it should be in the SHG, but it isn't. The OPO error signal also has no glitches despite being locked with ppol. It should also be noted that we didn't see any of these spikes in the fast or slow output monitor channels of either PLL.
Freezing one of the glitch spikes on the oscilloscope shows it has a width of 100 us, corresponding to a frequency of 10 kHz. This seems like a mechanical frequency, but at this point I had to leave for the day.

Yuhang, Michael
From Friday last week.
We recovered frequency independent squeezing to about 7 dB (vs 5 dB in April). The CC PLL is still glitchy and occasionally bumps the noise of the squeezed spectrum but the interval is enough to take an averaged spectrum.
CC PLL
Initially we attempted to check if the connection of the PLL actuation to the CC laser was broken. We saw that the slow signal monitor of the CC board has voltage, but the actuation cable from the slow servo did not.
After fixing this issue, we tried some other adjustments of the PLL parameters. Eventually we saw 10 MHz sideband detuning locks well. We went to lunch, came back and it was still locked.
We decided to check the sign behaviour of the PLL control loops. There are two control loops for each PLL, "fast" and "slow". The "slow" loop corrects long term drift. For each loop, we held the PLL sidebands at a location on the spectrum analyzer and plugged the output monitor to the oscilloscope. Then we saw the following:
Fast loop: When the fast loop is closed, the voltage of the correction signal goes up while the sideband frequency separation from the carrier is reduced. This implies a negative sign for the fast loop which is consistent with negative polarity set in the PLL software.
Slow loop: When the slow loop is closed, the correction signal goes down while the sideband frequency separation from the carrier is also reduced. So the slow loop has a positive sign. "Inv" should be off.
We were already using the smallest gain but the control loops still overshoot. We tried removing 12 dB attenuators, and this seemed to make the CC PLL lock good enough.
Frequency Independent Squeezing
We optimized ppol further by making the CC1 error signal large. We started from a ppol frequency of 160 MHz. But actually 160 MHz was already good enough and CC1 locked.
CC2 error signal was seen to go up and down on a timescale of 0.5s which is the correct behaviour.
Using the SR785, we saw -132.34 dBVrms/rtHz shot noise at high frequency (9 kHz) while blocking homodyne. CC2 fast loop seems to still be sending a lot of glitches to bump the squeezing spectrum every now and then, but the interval between glitches is still long enough to obtain 100 average traces. The presence of glitches is indicated by the reading of the green phase shifter high voltage driver going to zero, which indicates that the problem is somewhere in the CC signal. To speed up the inspection we just looked at very high frequency of 80 kHz. The lowest level of noise we saw was -139.2 dBvrms/rtHz, or 6.96 dB of squeezing.

Yuhang, Michael
From Thursday last week. We made many fixes to just about all parts of the frequency independent squeezer.
SHG
The RFPD that detects the SHG error signal was moved to the reflection of the first dichroic in the green path, aided by a lens and a green block filter. The error signal becomes like the reference on the wiki. We took the open loop transfer function eps1/eps2 on the spectrum analyzer but it was strange compared to the reference version (Aritomi thesis). We tried on the old spectrum analyzer and it seemed fine, with the correct low frequency response and first mechanical mode at 21 kHz. So maybe we just missed some settings. We changed the servo gain to give 2.2 kHz unity gain frequency and 45 degree phase margin. Coherence was close to one apart from a spike at 500 Hz.
OPO optimisation
The OPO was optimized using a fast method: scan green phase (ramp signal sent to green phase shifter high voltage driver) without locking OPO and measure amplification/deamplification of BAB. We saw that for some reason the green beam was greatly misaligned in pitch on both the filter cavity and OPO paths, so we adjusted with the first steering mirror after the green FI. Then green to OPO was maximized using the second 45 degree incidence mirror after GRMC transmission to maximize nonlinear gain. We saw:
74 mV without green
268 mV with 25 mW green, before optimization
496 mV after optimization
giving a nonlinear gain of 6.7. This is really quite high compared with what we saw before. We did not see the need to adjust temperature (though maybe we should).
IRMC
We had issues with mutual lock of SHG and IRMC ever since switching to the one EOM scheme. Currently SHG and IRMC are both modulated with one DDS channel and demodulated with one DDS channel. We took some cables to add phase delay to the IRMC and got the good PDH signal. So now it locks properly, although the threshold needs to be optimized a bit.
Homodyne alignment
The homodyne was realigned and balanced. The coarse alignment steps are as usual:
1) Block all OPO side infrared signals from reaching the homodyne
2) Lock IRMC to send local oscillator to homodyne
3) Adjust pitch and yaw of lens closer to IRMC to bring the homodyne photocurrent close to zero (will go far down if misaligned)
4) Adjust pitch and yaw of lens closer to edge of table to bring homodyne photocurrent close to zero (will go far upif misaligned)
5) Adjust balancing beam splitter to finish bringing the HD signal to zero
Then local oscillator is aligned to alignment mode cleaner for fine alignment.
The spectrum of the homodyne output was seen to be flat to about 40 Hz, with 100 Hz peak from ceiling light. We turned off the ceiling light but then it somehow turned back on. We decided to investigate using the security camera system but could not login.
After LO is aligned, the squeezed path is aligned by sending BAB to the alignment mode cleaner, but first we have to lock OPO with ppol.
GRMC
For whatever reason GRMC misaligned a lot as well. Maybe had something to do with the green misalignment in pitch above. During this process we could see the filter cavity reflection moving a lot on the green injection mirror though (there was an earthquake), so the green beam shouldn't be too far misaligned to FC.
PPol
The ppol frequency was changed to optimize BAB transmission (without green). It goes to 220 MHz, down from 260 MHz some time ago.
CC
On this day we saw that the CC PLL had erratic locking. The sidebands would come close to lock and then eventually quickly be ejected away from the resonance point.

Nishino,
By scaaning frequency, I measured the mode-mtching ratio of the current setup.There were 4 HOMs (both HG and LG). The mode-matching ratio is:
4.90 / (4.90 + 0.337 + 0.097 + 0.097 + 0.097)*100 = 88.7 %
The injection current was 1.1 A.
Also I locked the cavity by Moku:lab. Environmental sound seems to be too big and the locking was not stable only with a feedback onto PZT. It would need something like a shield for a stable locking.

Yuhang, Marc, Michael
We once again tried to find transmission sidebands at 88 MHz for the SHG locking but could not see them, so we had to rethink the lock scheme.
We had some discussion of how to resolve the BAB direct reflection issue (in fact, the beam is forming an interferometer with OPO HR and SHG input as the two arms, and SHG refl as the output port).
Our first thought was to put a Faraday isolator to decouple the two beams. For the SHG input path, there is not very much space for an FI, while on the BAB path the beam is quite large and would require a redesign of the mode matching telescope for BAB to OPO.
We had some thought about trying to decouple them using polarization. But given that SHG must be p-pol and BAB must be s-pol, we cannot decouple them using just HWPs and one PBS, and besides, we need those polarization optics for the Taiwan and Korea activities in ATC.
We then decided to try a pick-off from the SHG reflection before it reaches the SHG/BAB separation beam splitter and recombines with BAB. Since the current SHG refl PD has an ND2 attached and has about ~ 100 mW incident, we only really need about 1 mW of power to get a good signal. At this point we remembered that there are two dichroic mirrors on the green output of the SHG that reflect away residual IR into beam dumps. Actually, they reflect quite a lot, we measured 3.6 mW infrared reflected from the first dichroic after green generation.
And then we found that our attempts to see the transmission sidebands on the transmission PD were separately mistaken. On the first attempt, a low bandwidth PD was used in reflection, which is not fast enough to see RF sidebands. On the second attempt, an ND2 filter was left screwed on to the SHG refl RFPD when it was moved to the transmission side. So in both cases the sidebands were suppressed. Correctly implementing an RFPD in transmission of the SHG with sufficient incident power resulted in recovery of the error signal. So we didn't really have to change the locking scheme very much if at all.
We decided to put both SHG lock PDs in transmission. We selected a BSF10C beam sampler with an output wedge - this has ~ 1% reflection for p-pol at 45 degrees AOI, and maybe 5% at imperfect angle. To improve the visibility of the small transmitted power, the SHG correction signal SERVO OUT was disconnected, and an offset applied to the SHG PZT to see strong green generation. Then the IR transmission also becomes stronger, to about the level we would expect during SHG lock - about 3 mW transmission. In 99% transmission of the beam sampler I placed the RFPD Thorlabs PDA05CF2 which goes to the SHG error signal. In 1% reflection I placed the DCPD Thorlabs PDA36A-EC switchable gain, which goes to the servo threshold check and TRANSMIS OUT. I looked at TRANSMIS OUT on the oscilloscope -it was quite small so I increased the DCPD gain to 30 dB and it became about 550 mV. Then I checked the RFPD signal going to the mixer, which has 4.5 V max signal. I could see 95.7 % mode matching (4480 TEM00 + (96+56+48) HOMs). Then I placed a large beam dump at the former SHG refl PD point, which has quite large infrared power.
I set the threshold to slightly less than half of the DCPD signal (about 250 mV) and tried to lock, but the SHG would still only lock to a bad point (i.e. not to the level of the TEM00 peaks on TRANSMIS OUT), I looked at the error signal and tried to optimize the SHG demodulation phase but could not. From the reference signal on the wiki, it should have jump from a large negative point to a high positive point with a strong positive slope at the lock point (once again the NAOJ elog image upload issue is not very nice). However, instead it goes from large negative peak -> small positive peak -> small negative peak -> large positive peak, essentially making a small bump in the middle of the error signal with the opposite sign compared to what the polarity of the large peaks should give. I cannot get rid of this by changing the demodulation phase. Tomorrow we will try using the dichroic reflection for RFPD to see if that fixes the error signal.

Yuhang, Michael
We continued having SHG lock issues when investigating the stability of the green generation/stabilization path. We checked LO, it is at the correct frequency and level (5 dBm). The signal from the photodetector at the SHG mixer also looks fine on the spectrum analyzer (88 MHz peak going up and down with automated cavity scan). However, the output of the mixer PD x LO is quite bad. It was eventually found that direct reflection from BAB is coming back the the SHG reflection PD that we use to lock.
Originally, locking of the SHG was done via transmission through the SHG modulated at 15.2 MHz. This is within the linewidth of the SHG, so the PDH sidebands have strong transmission. However, it also was causing phase noise in squeezing due to being inside the linewidth of other green optics, so we decided to replace the 15.2 MHz modulation with 88.3 MHz. Since we expected the PDH sideband transmission through the SHG to be much weaker (and it is), we placed a PD in reflection of the SHG to lock. However, the reflection signal is transmitted through the same BS that splits the main laser to BAB.
Since I can't upload images right now I will just draw text art:
^
I
I
V
To and from BAB
\
<------- To SHG----- \ ---------> To SHG refl PD
\
^
I
I
From main laser
In fact, the BAB is injected at the OPO HR surface, so it is also reflected quite strongly, and the BS is 80:20 R:T, so there would be quite a lot of BAB reaching the SHG refl PD.
I tried to switch the error signal acquisition PD to the SHG transmission one - I took the cable connected to SHG refl PD (the one with the DC block) and connected it directly to SHG tra. I could see a nice cavity mode sweep on the oscilloscope, about 1.4 V or so, however, I could not see any mixer signal, nor could I see any 88 MHz sideband on the spectrum analyzer. There was a 0.2 ND filter going to the PD which I removed, but no difference. I tried to lock with very low positive, or negative, threshold on the servo but still nothing. So we have to reconsider how we lock SHG, otherwise we cannot have BAB and green at the same time, which is necessary to optimize the OPO for green and infrared.
I returned the ND filter put the DC block cable back to SHG refl.

Measurement finished and could find voltages range where both azimuth and ellipticity are crossing 0.
Then, with help of Yiru Zi-Hao and Byeong-Jun we moved the laser source to ATC clean room.
We did not brought back any replacement source yet.

Yuhang, Chang-Hee Kim, Michael
BAB to OPO - the BAB alignment to OPO was optimized using the hom sub dc signal with one homodyne input blocked. Mode matching was increased to 1170 mV + (46, 34, 14, 18 mV) -> 91.8%. Then ppol was also aligned from a mode mismatch of about 50% and now OPO locks with ppol. However, since the OPO temperature hasn't been optimized for a long time we should re-optimize parametric gain.
PLL - there is still quite a lot of signal in the ppol PLL so probably it was not misaligned too much to the fiber (I accidentally misaligned it in pitch last time when I thought I was adjusting BAB to OPO).
Lasers - We adjusted the temperature and current for each laser to be consistent with the values in the wiki +/- 0.001
GRMC MZ - we want to measure OPO nonlinear gain by injecting green and measuring BAB transmission, but GRMC/MZ kept unlocking on short timescales for some reason. The offset for the MZ HVD was unintentionally left on by me but it should be off. But the system kept unlocking anyway. So we decided in the end to take transfer function measurement, but it was getting late, so we will do tomorrow

Measurement was finished at 23deg. I rotated the mirror back to 0deg. There was some discrepancy in measurement at this position. The measurement is ongoing in between 1.34-1.36 and 16.42-24.62 with 1mV resolution for first and second LC respectively.
foldername=r'C:\Users\atama\OneDrive\LC-Experiment\Measurement Data\Coating measurement\BB1-E03\0 deg'
filename='Mon, Dec 18, 2023 9-16-57 AM.txt'

I designed 3 parts in SolidWorks for testing the hydrogel birefringence change vs applied force.
1 holder for the DM713 use to apply force, 1 cube to apply uniform force and 1 holder for the cell container.
Files are saved in my folder of the workstation in room 216/217.
Hirata-san is helping for the 3d printing that will be done at ATC.

I took a Thorlabs 1064 VLP Faraday isolator (4.7mm aperture, max power 1.7 W) from the speed meter experiment to the KOACH clean room. We still need an appropriate power meter - we would like to inject up to 1 W into the SHG.

The voltage scan was performed for the measurement. The scans started from 0-25V for each LC with a resolution of 0.1V and were finished with 1-1.3V and 22-23V with 0.001V resolution for the 1st and 2nd LC respectively.
I have rotated the mirror by 23 deg. The voltage scans were started for 0-25V range with 0.01V 0.1V (edited) resolution.
measurement file is here:
C:\Users\atama\OneDrive\LC-Experiment\Measurement Data\Coating measurement\BB1-E03\23 deg\Wed, Dec 13, 2023 1-20-43 PM.txt

MZ and GRMC
Lock was recovered with the nominal level of stabilized output power (power meter in transmission of GRMC says 26.5 mW green to OPO at 4.2 on the MZ offset dial, which is our usual setting).
I tried mode matching on each arm -
PZT arm: 384 mV TEM00, 8 mV HOM, 96.9%
Non PZT arm: 364 mV TEMO00, 33.2 mV total HOMs, 91.6 %
After adjusting them together, then maximising the mode cleaner transmission using the MZ HVD offset (different control knob to the above), I get:
GRMC: 1.5 V TEM00, 68 mV HOMs, 95.76%
but one other time I got 720 mV, 72.4 HOMs, 90.9%
The power meter autoranging should be off. Maybe I am just a bit confused about the MZ scan. A bit of explanation for people not familiar, the MZ is not locked using PDH since it is not a cavity, rather it is just locked to some reference voltage controlled by MZ offset dial. When the LOCK switches are turned on, the GRMC scans across the free spectral range quickly and acquires lock first using the PDH sideband signal from GRMC reflection PD, then the MZ scans slowly and locks second using the signal from GRMC transmission PD to keep the transmitted green power at a stable level. Also, the interference condition of MZ using the HVD offset can be changed to reflect or transmit TEM00/HOM - we want it to transmit mostly TEM00 and reflect HOM into a rejected port. Anyway, I don't remember how slow is "slow" scan, but I don't think it should be causing the slow change of MZ mode matching on the timescale I saw. Either way, that is only an issue for measuring mode matching when not locked. The system locks and sends stabilized power to OPO so we can fix it later.
BAB to OPO
I recovered bright alignment beam mode matching to about the level it was during Yuhang's last visit, which is to say, ok but not great (~ 60%). I measured using power meter in transmission of OPO - it is quite hard to see the BAB transmission with OPO unlocked since it is injected from the HR side of the OPO, but there was a conveniently placed holder at just the right spot for the power meter. I tried to measure with homodyne but the IRMC started having some issue where it wouldn't lock to the correct point. We have been having issue with SHG and IRMC mutual lock ever since we switched to the single 88 MHz EOM to do all of the squeezer table modulations. We should decouple these two signals some time.
FC green lock
I turned on the FC length control servo inside the clean room and then tried to lock FC on green but was not successful.

I could fix 2 out 3 metric screws in the KAGRA mir box (spareETMY). the last screw was harder to fix and I didn't had the proper key so I did not fix it yet..
The spare screws are in the PCI pre-clean room

The IRMC was not too badly misaligned, just tweaking a bit the lens in front of it I could recover about 94% mode-matching.
Further improvement would require to tune the IRPS but it's not high priority for now so I left it like that.
Then, could smoothly lock it.

[ByeongJun, Marc, Michael]
The GR beam was realigned on BS chamber using PR picomotors, on 1st and 2nd target using BS picomotors.
We tuned END coils actuators to get the beam on the back of 2nd target.
Then, using IN picomotors, we recovered fundamental mode in transmission of FC.
To do :
offload END pitch (now 280 is sent to coil)
replace the quad box for the camera display close to DGS
recover remote control of 2nd target

I replaced the newfocus 5104 with Thorlabs BB1-E03 mirror.
Actually, when I first checked the beam centering, I noticed that the beam was not centered on the 5104 mirror. The beam at this instance was at the center of the camera.
So, even though I changed the mirror, it was made sure that the beam was incident at the center of mirror, and the reflection was at the center of the camera. In, this process only the mirror was moved. All other components were left undisturbed.
The LC voltage scan was done from 0-25V for both with 0.02 resolution.
measuerment data location
C:\Users\atama\OneDrive\LC-Experiment\Measurement Data\Coating measurement\BB1-E03-10\Fri, Dec 8, 2023 5-48-31 PM.txt
Edited:
measurement file is here:
C:\Users\atama\OneDrive\LC-Experiment\Measurement Data\Coating measurement\BB1-E03\0 deg

I confirmed the alignment of the green AOM.
The (lower left on electronics rack) function generator outputs from channel 1: 109.036 035 615 MHz, 5 dBm voltage, to + 18 dB RF amp -> 23 dBm to AOM.
I reduced the main laser current from 1.833 to 1.750 to make the green power consistent with previous values of AOM green power. I confirmed the alignment of the first order diffracted beam from the AOM to the subsequent iris and the green is now hitting the PR tank targets.
SHG green out: 284 mW
AOM in: 48.8 mW
AOM out (RF on): 46.0 mW
AOM out (RF off): 45.2 mW
Iris out (RF on): 22.1 mW
Iris out (RF off): < 10 uW
Measurement finished and could find voltages range where both azimuth and ellipticity are crossing 0.
Then, with help of Yiru Zi-Hao and Byeong-Jun we moved the laser source to ATC clean room.
We did not brought back any replacement source yet.